Approximately how many grams of ammonium chloride should be dissolved in a liter of 0.3 F NH3 to reduce the concentration of hydroxide ions to one one hundredth of its original value.




1
Expert's answer
2021-03-29T05:15:04-0400

Molar Mass of NH4Cl = 53.491

Molar Mass of NH3 = 17.031

NH3 + HCl -> NH4Cl

Mole ratio = 1 NH3 = 1 NH4Cl

0.3 = N/ 17.031

N = 17.031 × 0.3 = 5.1093

5.1093/(53.491 + 17.031) = 5.1093/70.522 = 0.0724

0.0724 = 1/100x

0.0724 = 0.01x

x = 0.0724/0.01

= 7.24g of Ammonium Chloride